1. A researcher conducts an experiment comparing two treatment conditions and obtains data from 15 people for each treatment condition

a. If the researcher used an independent-measures design, how many subjects altogether participated in the experiment?

b. If the researcher used a repeated-measures design, how many subjects altogether participated in the experiment?

c. If the researcher used a matched-subjects design, how many subjects altogether participated in the experiment?
Answer #1

a) since there are two groups

total number of subjects required =2*number of subject in each group =2*15 =30

b)

since same subjects are in both groups

total number of subjects required =15

c)

since subjects are different in both groups( only they are matched with some attribute):

total number of subjects required =2*number of subject in each group =2*15 =30